# Are Economists Always More Introverted? Analyzing Consistency in Persona-Assigned LLMs

### v1

@@ 5 Related work
Persona-assigned LLMs. Personas can guide LLMs to generate responses that align with specific values and beliefs Li et al. (2024); Santurkar et al. (2023). However, they can also expose stereotypes embedded in the model Park et al. (2024); Gupta et al. (2023), raising concerns about bias and unintended implications. Persona adherence is usually evaluated using self-report scales, but Wang et al. (2024b) use interview-based testing to capture actual model behavior, showing the need for application-specific evaluations. Malik et al. (2024) examine how different personas from various sociodemographic groups influence writing styles. Apart from inference-time persona assignment, it is also possible to further fine-tune the model. For example, Shao et al. (2023) train LLMs to adopt specific personas using three key components: a profile (a detailed persona description), a scene (a situational context), and interactions. Wang et al. (2024a) suggest including dialogues for persona assignment of LLMs. Our framework can evaluate these realistic personas or finetuned models that do not perfectly fit our predefined categories, as shown in Appendix H.
@@ 8 Ethical considerations
Appendix A Persona Categories
@@
Appendix B Persona prompts
@@ C.2 LLM-as-a-judge
Appendix D Model Checkpoints
@@
Appendix E Characteristic-specific consistency
@@
Appendix F Model Comparisons
In this section we focus on the following research question: Do consistency patterns differ across model families and/or within a single model family?
@@ Appendix F Model Comparisons
Figure 8 illustrates how consistency varies across model families. For example, we see that Ministral-8B has lower overall consistency than Llama-3.1-8B despite similar model size. Additionally, our results show that within a model family, larger models tend to be more consistent than smaller models. This is shown by the three Llama models in the figure. This result aligns with the finding from Serapio-García et al. (2023), showing higher reliability and validity of synthetic LLM personality for larger and instruction fine-tuned models. Additional statistical analysis is reported in Appendix G.
Appendix G Statistical Analysis
Additional statistical analysis was done on the results for the model and dimension comparisons.We applied a one-sided Wilcoxon signed rank test with a confidence of 95%. For all models the intra-persona entropy is significantly lower than the inter-persona consistency (all p-values <0.01). Additionally, when applying the same test to the different dimensions, we find no statistical differences for the singlechat and multichat evaluation dimensions (p-values are 0.1012 and 0.0948 respectively) for the other dimensions we find statistical significant differences (p-values < 0.001).
@@ Appendix G Statistical Analysis
Llama 3B 1.70 1.70
Llama 8B 3.50 2.65
Llama 70B 3.55 3.85
Ministral 2.45 2.55
Qwen 3.80 4.25
@@
Survey 3.50 3.85
Essay 3.45 3.15
Social Media 4.10 4.25
Singlechat 1.55 1.20
Multichat 2.40 2.55
@@
Furthermore, to identify a ranking among the different experiments, we conducted a Friedman test to identify whether there are significant differences and followed this with a Nemenyi test, given that our result showed there are significant differences to include a ranking. Tables 2 and 3 show the rankings of the different models and dimensions. Qwen is in terms of inter-persona consistency only statistically not different from Llama 70B and on the intra-persona consistency, only Llama 3B is statistically different both on a 95% confidence interval. For the different dimensions, we find statistical differences between survey, essay, and socialmedia on one hand and both singlechat and multichat for the intra-persona consistency on the other hand. For the inter-persona consistency, survey, essay, and socialmedia are significantly outperforming singlechat and only socialmedia post generation is significantly outperforming multichat.
Appendix H Real-world Applicability
A last and additional research question we aim to investigate is whether our framework can be applied in realistic settings where personas do not perfectly align with predefined categories. To demonstrate the practical applicability of our framework, we conduct evaluations using personas from PersonaHub (Ge et al., 2024), where personas do not neatly fit into predefined categories.
We assess consistency in a realistic scenario for Qwen to illustrate the real-world applicability of our framework using five randomly selected personas from the Personahub from  Ge et al. (2024). We chose Qwen as it provides representative results for all models with an average correlation of 0.70 on the first experiments. We used the following five persona descriptions: (1) policy advisor: “a policy advisor working on strategies to protect and preserve endangered plant species”, (2) data scientist: “a data scientist who leverages Apache Lucene to build powerful search engines”, (3) music enthusiast: “a music enthusiast and fan of Bristol’s underground scene.”, (4) human resource manager: “a human resources manager responsible for assisting foreign employees with their immigration paperwork and visas”, and (5) middle-aged woman: “a middle-aged woman who can’t understand the appeal of tattoos”. We analyze the results in what follows

### v2

@@ 5 Related work
Persona-assigned LLMs. Personas can guide LLMs to generate responses that align with specific values and beliefs Li et al. (2024); Santurkar et al. (2023). However, they can also expose stereotypes embedded in the model Park et al. (2024); Gupta et al. (2023), raising concerns about bias and unintended implications. Persona adherence is usually evaluated using self-report scales, but Wang et al. (2024b) use interview-based testing to capture actual model behavior, showing the need for application-specific evaluations. Malik et al. (2024) examine how different personas from various sociodemographic groups influence writing styles. Apart from inference-time persona assignment, it is also possible to further fine-tune the model. For example, Shao et al. (2023) train LLMs to adopt specific personas using three key components: a profile (a detailed persona description), a scene (a situational context), and interactions. Wang et al. (2024a) suggest including dialogues for persona assignment of LLMs. Our framework can evaluate these realistic personas or finetuned models that do not perfectly fit our predefined categories, as shown in Appendix J.
@@ 8 Ethical considerations
Acknowledgements
This research was funded by the Statistics Flanders research cooperation agreement on Data Science for Official Statistics. The resources and services used in this work were provided by the VSC (Flemish Supercomputer Center). We acknowledge the support of the Research Foundation Flanders (FWO), Grant G0G2721N.
Appendix A Surveys
Below additional information regarding the surveys is included.
Happiness survey:
The happiness survey includes a survey of four questions. For each question, the participants have to indicate the point of the scale that they find most appropriate for them. An example question is: "In general, I consider myself: 1 (not a very happy person) - 7 (a very happy person)."
Political Stance:
This survey includes 62 statements where participants have to indicate on a four-point likert scale if they disagree or agree with the statement. The survey itself includes predefined scores to determine the quadrant to which the participant belongs. An example of a statement is: “If economic globalisation is inevitable, it should primarily serve humanity rather than the interests of trans-national corporations.”
Occupations:
In this survey, participants are asked to indicate what activities they like to do (e.g., fix electrical things, take art courses, attend conferences…), which competencies they have (e.g., I can repair furniture, I can do interpretive reading, I am a good debater…), what jobs they would like to do (e.g., machinist, musician, hotel manager…), and how they would estimate their own skills ( e.g., friendliness, clerical ability, managerial skills…).
Personality:
In the Big Five Personality test, participants are asked to indicate to what extent they think a certain characteristic might apply to them or not. Participants are asked to respond on a five point Likert scale. An example characteristic is: "Is helpful and not selfish with others".
Appendix B Persona Categories
@@
Appendix C Persona prompts
@@ D.2 LLM-as-a-judge
You **must** choose between one of the options. First output this category, then generate a confidence score, where 1 is Very unconfident, 2 is unconfident, 3 is confident, 4 is very confident. Use the following structured output and fill in [category] and [confidence_score]: choice:[category],confidence:[confidence_score]
Appendix E Entropy
Entropy is a well-suited metric for consistency analysis, as it quantifies the uncertainty in the distribution of category assignments for each persona category (e.g., occupation, happiness, personality, and political stance). We chose this metric because of the following reasons:
  • Applicability across different persona categories, accommodating both binary and multiclass traits.
  • Capturing full distributional patterns, considering all assigned categories rather than just the majority class.
  • Enabling comparability across personas and models without arbitrary thresholds.
In literature, this metric is used to assess consistency (e.g., Lyu et al. (2025)). However, entropy does not reveal which specific label a model is consistently assigning. To address this, we include the characteristic-specific consistency scores as well in our analysis.
Appendix F Model Checkpoints
@@
Appendix G Characteristic-specific consistency
@@
Appendix H Model Comparisons
In this section we focus on the following research question (RQ4): Do consistency patterns differ across model families and/or within a single model family?
@@ Appendix H Model Comparisons
Figure 8 illustrates how consistency varies across model families. For example, we see that Ministral-8B has lower overall consistency than Llama-3.1-8B despite similar model size. Additionally, our results show that within a model family, larger models tend to be more consistent than smaller models. This is shown by the three Llama models in the figure. This result aligns with the finding from Serapio-García et al. (2023), showing higher reliability and validity of synthetic LLM personality for larger and instruction fine-tuned models. Additional statistical analysis is reported in Appendix I.
Appendix I Statistical Analysis
Additional statistical analysis was done on the results for the model and dimension comparisons.We applied a one-sided Wilcoxon signed rank test with a confidence of 95%. For all models the intra-persona entropy is significantly lower than the inter-persona consistency (all p-values <0.05). Additionally, when applying the same test to the different dimensions, we find no statistical differences for the singlechat and multichat evaluation dimensions (p-values are 0.1012 and 0.0715 respectively) for the other dimensions we find statistical significant differences (p-values < 0.001).
@@ Appendix I Statistical Analysis
Llama 3B 1.80 2.05
Llama 8B 3.40 2.25
Llama 70B 3.65 4.10
Ministral 2.45 2.40
Qwen 3.70 4.20
@@
Survey 3.55 3.90
Essay 3.20 3.35
Social Media 4.20 4.05
Singlechat 1.60 1.15
Multichat 2.45 2.55
@@
Furthermore, to identify a ranking among the different experiments, we conducted a Friedman test to identify whether there are significant differences and followed this with a Nemenyi test, given that our result showed there are significant differences to include a ranking. Tables 2 and 3 show the rankings of the different models and dimensions. Qwen is in terms of inter-persona consistency only statistically not different from Llama 70B and on the intra-persona consistency, only Llama 3B is statistically different both on a 95% confidence interval. For the different dimensions, we find statistical differences between survey, essay, and socialmedia on one hand and both singlechat for the intra-persona consistency on the other hand, as well as between survey and socialmedia on the one hand and multichat on the other hand. For the inter-persona consistency, survey, essay, socialmedia, and multichat are significantly outperforming singlechat and only socialmedia post generation is significantly outperforming multichat.
Appendix J Real-world Applicability
A last and additional research question we aim to investigate is whether our framework can be applied in realistic settings where personas do not perfectly align with predefined categories (RQ5). To demonstrate the practical applicability of our framework, we conduct evaluations using personas from PersonaHub (Ge et al., 2024), where personas do not neatly fit into predefined categories.
We assess consistency in a realistic scenario for Qwen to illustrate the real-world applicability of our framework using five randomly selected personas from the Personahub from  Ge et al. (2024). We chose Qwen as it provides representative results for all models with an average correlation of 0.66 on the first experiments. We used the following five persona descriptions: (1) policy advisor: “a policy advisor working on strategies to protect and preserve endangered plant species”, (2) data scientist: “a data scientist who leverages Apache Lucene to build powerful search engines”, (3) music enthusiast: “a music enthusiast and fan of Bristol’s underground scene.”, (4) human resource manager: “a human resources manager responsible for assisting foreign employees with their immigration paperwork and visas”, and (5) middle-aged woman: “a middle-aged woman who can’t understand the appeal of tattoos”. We analyze the results in what follows
@@
Appendix K Combination of different persona categories
Exploring “in-between” personas,where combinations of these categories are specified, would provide valuable insights into the interplay between traits in terms of persona-assigned consistency. Given the high computational expenses to do a full exploration of the different combination, we already provide some initial insights by conducting additional experiments combining personas across categories. More specifically, we experimented with the combination of happy/sad and pilot/economist, as shown in Table 5. These initial results indicate that combining occupation and happiness personas primarily increases consistency within those categories, with some spillover effects observed in personality consistency. However, the political category remains challenging to express consistently with this combination of personas.
Personas Evaluation Categories
Happiness Occupation Personality Political
A happy economist $0.08\pm 0.17$ $0.13\pm 0.18$ $0.22\pm 0.08$ $0.76\pm 0.25$
A happy pilot $0.0\pm 0.0$ $0.31\pm 0.30$ $0.26\pm 0.19$ $0.94\pm 0.08$
A sad economist $0.17\pm 0.24$ $0.22\pm 0.24$ $0.45\pm 0.13$ $0.85\pm 0.22$
A sad pilot $0.06\pm 0.13$ $0.23\pm 0.31$ $0.48\pm 0.12$ $0.80\pm 0.44$